OpenStudy (anonymous):

What are the coordinates of the midpoint between the points (4, 5) and (12, 5)? A. (5, 5) B. (6, 5) C. (7, 5) D. (8, 5)

OpenStudy (anonymous):

is there a picture?

OpenStudy (anonymous):

No, no picture. :(

OpenStudy (anonymous):

again ... what number is halfway between 4 and 12? what number is halfway between 5 and 5?

OpenStudy (anonymous):

maybe this will help

OpenStudy (anonymous):

I have no idea... :(

OpenStudy (jack1):

@guest.002 is correct... coordinate is (x, y) x = what number is halfway between 4 and 12? y = 5, as both y values are 5 equation is (x1 + x2)/2

OpenStudy (jack1):

and (y1+y2)/2

OpenStudy (anonymous):

Would the answer be B? @Jack1

OpenStudy (jack1):

no (x1, y1) and (x2,y2) (4, 5) and (12, 5) our point =(x3,y3) x3 equation is (x1 + x2)/2 = (4 + 12) / 2 = 16/2 = 8 so x3 = 8 y3 = (y1+y2)/2 = (5 + 5)/2 = 10/2 = 5 so point is ( 8,5)

OpenStudy (jack1):

does that make sense @rebecca007 ?

OpenStudy (anonymous):

Thank you SO much. I get it now!!

OpenStudy (jack1):

sweet as, yw, slaters